Staff:

LBTO: Juan Carlos

Observer: SHARK-NIR team (remote)

Details.

Chamber temperature : -3.25C

08:30
  • WFS up
  • OCAM
  • ok
  • Argos cal unit.
  • ok.
  • flat:
  • 15:49:44.742018| MAIN > Maximum mode index used to apply flat -> # 639
test preset. Acquiring PSF

Thu Mar 2 16:07:26.731 2023 aos.info - WFS source set to: LBTIWFS
Thu Mar 2 16:07:27.478 2023 aos.info - FLAO: WFS: loading setup for SHARK
Thu Mar 2 16:07:31.993 2023 aos.info - FLAO: WFS: configuring CCDs
Thu Mar 2 16:10:38.419 2023 aos.PresetAO{FLAO}.failed - PresetAO {FLAO}[10] - [AOException] presetAO: FATAL - Reading shared buffer mirrorctrl.L:TECHVIEWER_65536 (-5001) TIMEOUT_ERROR
  • The CCD47 not live taking a dark
  • BCU47 stuck in the taking a dark for the CCD47
  • power cycle bcu47
  • ok.
  • configure the ccd47 from the HW gui
  • OK
  • manual dark from the wfs control gui
  • presetao
  • ccd47 not live.
  • reset BCU57 from PDU
  • configure ccd47
  • ok
  • take a dark
  • ok
  • presetao
  • ok
  • ?
  • To be investigated
  • restart the auxloop.
  • not display of AO telemetry
  • sterDiagnostic_28904|ERR| 61|2023-03-02 16:39:55.008328| OPTICAL-LOOP > Rejected wrong filotto between WFS and AdSec: 00000000 and 001CCA81
    MasterDiagnostic _28904|ERR| 62|2023-03-02 17:48:53.701708| OPTICAL-LOOP > Rejected wrong filotto between WFS and AdSec: 00559FE6 and 00000000
    MasterDiagnostic _28904|ERR| 63|2023-03-02 17:51:36.593356| OPTICAL-LOOP > Rejected wrong filotto between WFS and AdSec: 0057DC88 and 00000000
    MasterDiagnostic _28904|ERR| 64|2023-03-02 18:05:33.494370| OPTICAL-LOOP > Rejected wrong filotto between WFS and AdSec: 00000000 and 00635A9B
09:38
  • loop closed.
  • TN:20230302_163544
  • SHARK running.
09:44

TN:_164431

10:00
  • adjusting PSF in shark camera, moving baysisde stage x and y
  • x->66.5, y->21.9, z->62
  • TN:_170043
10:487
  • shark team working on.
  • TN:_174552
10:55
  • move z 59
  • TN:_155505
11:07
  • move z to 62
  • TN:_180656
  • TN:_181550
11:22
  • move z to 59
  • TN:_182803
11:35
  • move to 62
  • TN:_183448
12:01
  • move z to 59
  • TN:_190048
12:07
  • move z to 62
  • TN:_190705
12:13
  • improved the vibrations
  • total jitter 6mas.
  • TN:_191311
12:40
  • move z to 59
  • TN:_104027
12:45
  • move to z 62
  • TN:_194531
1:17

move z to 59

TN:_201729

1:28

oscillation in the pipoils.

dome rotating.

1:38
  • resume close loop
  • move to 62
  • TN:_203839
1:53
  • enclosure rotation.
  • pause loop. gain zero
2:24
  • resume loop
  • TN:_213426
  • TN:_224259
4:00

Stop.

Argos parked.

WFS power off, pdu; 5,8,15,16 off

-- JuanCarlosGuerra - 02 Mar 2023
Topic revision: r1 - 02 Mar 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback